OpenStudy (anonymous):

Pretty Please? What are the oxidation numbers for each element in: Fe + O2 ---> Fe2O3 ? Can you please show me how to do the problem?

OpenStudy (anonymous):

@dtan5457 @aaronq @abb0t @cp9454

OpenStudy (chmvijay):

oxidation state of element in its elemental form is 0 oxidation state O is -2

OpenStudy (aaronq):

You start off with a basic set of rules and you figure the rest out with simple arithmetic. Rules (most often true, except in rare occasions): Group 1 always +1 Group 2 always +2 Oxygen always -2 Group 7 always -1 Noble gases, elemental or homonuclear (the same element) molecules are zero The rest you have to figure out by subtracting or adding to achieve whatever charge the compound bears. eg. H2O H is group 1, so H = +1, there are 2 of them O must balance the charge so, O=-2 HCl H=+1, Cl=-1 MgSO4, you have to know that SO4^2- is a polyatomic ion, treat that as it's own separate thing: O=-2 Total contribution of O=-2(4) = -8 ^4 oxygen atoms Charge on ion=-2, so S has to fulfil the rest which is 6, so S=+6 back to MgSO4, so we treat SO4 as entity and it bears a -2 charge, Mg therefore has to be +2 to balance the charges.

OpenStudy (aaronq):

you can also look up tutorials for this on youtube

OpenStudy (aaronq):

if you wanna practice, try some, i'll check them NH4+ NO2 N2O PCl3 CuBr2

OpenStudy (anonymous):

Haha, thanks Aaron. Unfortunately i have an entire sheet of them to do for homework. You rock. :) Are you a chem major by any chance?